OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

cos 3x is cos (2x + x) so you apply cos (A + B) and cos(2x) here

OpenStudy (anonymous):

do you know what those are?

OpenStudy (anonymous):

Yes those are familiar. cos a cos b + sin a sin b = something right?

OpenStudy (anonymous):

wait it equals the cos(a+b) yeah

OpenStudy (anonymous):

cos (A+B) is cosAcosB - sinAsinB

OpenStudy (anonymous):

so cos(2x+x) = cos 2x cos x - sin 2x sin x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then cos 2x is cos^2 x - sin^2 x

OpenStudy (anonymous):

if you want to see why cos (2x) = cos (x + x) = cosx cosx - sin x sin x thus cos^2 x - sin ^2 x

OpenStudy (anonymous):

do you know the value of sin(2x)?

OpenStudy (anonymous):

sin(2x) would be sinx sinx - cosx cosx ?

OpenStudy (anonymous):

nope. sin(A+B) is sinAcosB + sinBcosA so sin(2x) = sin(x+x) = sinxcosx + sinxcosx

OpenStudy (anonymous):

that gives you 2sinxcosx

OpenStudy (anonymous):

a)cos x - 4 cos x sin2x b) -sin3x + 2 sin x cos x c) -sin2x + 2 sin x cos x d) 2 sin2x cos x - 2 sin x cos x Those are the answers. I think all we need to do now is simplify

OpenStudy (anonymous):

yes simplify everything